One of two supplementary angles is five times the other. Find the angles.
(Recall that supplementary angles are angles whose sum is equal to 180 .)

Answer by bucky(2189)

Let the smaller angle be represented by x. The bigger angle is 5 times the smaller,
so the bigger angle is 5*x.
.
When you add the two angles together, the sum is 180 degrees.
.
In equation form this is:
.
x + 5x = 180
.
Adding the two terms on the left side results in:
.
6x = 180
.
To solve for x, divide both sides by 6 to get:
.
x = 180/6 = 30
.
Since x is the smaller of the two angles, we know that the smaller angle is 30 degrees.
And the larger angle is 5 times as big so it is 5*30 = 150 degrees.
.
As a check ... does 30 + 150 = 180 degrees? Yes it does and so our answer is correct.
.
Hope this helps you to understand the problem a little better.